CRY+VIS Interface Meeting
Date: 2021/12/3 13:00-13:45
Participants: Takahashi, Yasui, Sugiyama, Yokozawa, Ikeda, Tomaru, Hirata, Miyoki, Uchiyama, Sato, Ushiba, Akutsu
- IX: Finalization of cryo-payload. Replacement of flip cables on 12/2. Cable repair on 12/6.
- EX: Finalization of cryo-payload. Height adjustment of TM on 11/30-12/1. Cleaning of the cross tube on 12/7.
- EY: There are some problems (BF actuators, F0 yaw FR). Will be checked on 12/7.
- Replacement of flip cables: IX on 12/2.
- TM height setting: ITMY.
- Oplev: Setup of the TM Oplev in IY after the TM alignment.
- ACC: Driver setup for IY on 12/2. Assembly for EY is ongoing.
- Commissioning in air: Return of GRY.
- Commissioning in vacuum: Chamber closing is scheduled from 12/13.
- Overall commissioning:
- Acceptance check: Review by 12/8.
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: IXV and EXC-BRT on 12/9-10.
- Heater control:
- Power source for Oplev: Finished in IX.
Date: 2021/11/26 14:00-14:15
Participants: Takahashi, Hirata, Sugiyama, Ishizaki, Terrence, Tanaka
- Adjusted the height of TM in EX. The height of the BF damper and the F1 damper was adjusted as well.
- Replaced the flip cables in EX.
- Setup the drivers for the accelerometers in IY.
To try to eliminate the wandering noise in the BS F1 LVDT, we connected the BS chamber to electrical ground using a power cable: it did not have any effect (klog 19034).
- Before connecting, we measured the voltage betwwen the chamber and the ground: 17 mV rms.
- We also checked the resistance between the connector that plugged into the rack and the chamber: it was only 0.2 Ohm, indicating there wasn't any problem with the cable.
The wandering peak is not observed in any other GAS filters in Type-B suspensions (klog 19027).
- More ideas are welcome.
We calibrated the BS F0 LVDT. The new calibration factor and offset are already in the medm screen and accepted in the SDF (klog 19057).
- We changed flip adapters in the BS chamber.
- Hirata-san finished the mechanical design of the holder for the laser displacement sensor, which will be used to calibrate the BS F1 filter.
- While I found during the work:
- The IP is very soft, when applying DC offset to it, it doesn't return to its original position. And its position is always shifting. Is this an indication of it being close to unstable?
- The IP diagonalization (actuation at least) is not good. I will try to fix it next week.
- I measured transfer functions of the BS but something's wrong with the F1 GAS and IP T measurements. I will try to measure them again next week.
- I had been trying to make simple damping controls for SR2 but I haven't made much progress because I am trying to finalize the code that can automate this process.
- Type-A team: Review of IX and EX. Recovery of actuators in EY.
- Type-B team:
[Terrence] Fabian and I would like to lock the IP and F0 of the suspensions (BS, SR2, SR3) and measure the readouts for an estimation of the sensor noise. This information will be useful for properly designing blending filters in the future. We would like to do this when it's still possible to do so. Fabian suggested doing this next week.
(Date in Kamioka)
The week from 12/6
- Takahashi: 12/7-10
- Sato: 12/6-9
- Hirata: 12/7-10
The week from 12/13
- Takahashi: 12/13-17
- Sato: 12/13-17
- Hirata: 12/14-17